# MultiverSeg: Scalable Interactive Segmentation of Biomedical Imaging Datasets with In-Context Guidance

### v1

@@ Abstract
Medical researchers and clinicians often need to perform novel segmentation tasks on a set of related images. Existing methods for segmenting a new dataset are either interactive, requiring substantial human effort for each image, or require an existing set of manually labeled images.
@@
We demonstrate that MultiverSeg enables users to interactively segment new datasets efficiently, by amortizing the number of interactions per image to achieve an accurate segmentation. Compared to using a state-of-the-art interactive segmentation method, using MultiverSeg reduced the total number of scribble steps by 53% and clicks by 36% to achieve 90% Dice on sets of images from unseen tasks. We release code and model weights at https://multiverseg.csail.mit.edu.
Figure 1: MultiverSeg enables users to rapidly segment new datasets. The MultiverSeg network takes as input an image to segment and user interactions, along with a context set of previously segmented image-segmentation pairs (left). As the user completes more segmentations, those images and segmentations become additional inputs to the model, populating the context set. As the context set of labeled images grows, the number of interactions required to achieve an accurate segmentation decreases (right).
@@ 1 Introduction
Segmentation is an important step in biomedical image analysis pipelines for both biomedical research and clinical care. Biomedical researchers and clinical researchers often acquire novel images types or identify new regions of interest, and need to perform new segmentation tasks. Typically, scientists want to segment the same region of interest in many similar images from a new dataset.
Manually segmenting images is labor-intensive and requires domain expertise. Interactive segmentation systems, in which a user provides a few clicks or scribbles on an image to produce a predicted segmentation, help to speed up the annotation of individual images. But with existing interactive segmentation systems, the user must independently repeat the same process for each image [59, 126, 85, 82, 133]. It would be impactful if the segmentation system became more accurate as the user completed more segmentations from the same task, learning from experience.
We propose MultiverSeg, a new interactive system that progressively reduces the number of user interactions needed to predict accurate segmentations, as more images are segmented in a particular task (Fig. 1). MultiverSeg takes as input user interactions for a new image, along with a context set, of example (previously segmented) image-segmentation pairs. To segment a new dataset, the user begins by interactively segmenting the first image. Once completed, the example becomes an input to MultiverSeg, providing context for the segmentation of subsequent examples. The user interactively segments the next image with MultiverSeg using bounding boxes, clicks or scribbles. As the user labels more images, the context set grows and the number of interactions required to achieve the desired segmentation of the subsequent image decreases, often to zero. Unlike existing interactive segmentations systems [126, 59, 82, 85, 136, 123, 76, 77, 131, 107, 122] where the work required to segment a dataset is linear in the number of images, MultiverSeg enables users to rapidly segment entire datasets.
@@ 2 Related Work
Interactive Segmentation. Recent interactive segmentation models can generalize to new segmentation tasks in medical [126] and natural images [59, 103]. While these methods are effective for segmenting single images, they require prohibitively extensive human interaction when segmenting large datasets. To incorporate new information, they must be retrained or fine-tuned. In contrast, at inference time MultiverSeg can be conditioned on example segmentations from the new task, dramatically reducing the human effort required for accurate segmentation.
@@
A few works have explored in-context segmentation using a context set of example images with user annotations on those example images. OnePrompt [128] segments a medical image given exactly one context example with click, scribble, bounding box or mask annotation. In contrast, MultiverSeg has mechanisms to facilitate a variable number of context set entries and enables users to incorporate interactions on the target image even when given the context of previous segmentations, enabling substantially richer use cases. LabelAnything [23] is a few-shot framework that enables multi-label segmentation of a target natural image with a small variable size context set of example images with click, box, or mask annotations. In contrast, MultiverSeg can leverage large context sets, and enables users to provide interactions on the target image and make corrections to refine the prediction to get the desired segmentation.
@@
Annotation-Efficient Learning. Another approach to segmenting a new dataset is to collect sparse annotations on many images and train an automatic segmentation model using these annotations for supervision. The annotations may be bounding boxes [65, 117], clicks [75, 78, 106] or scribbles [71, 70, 135, 83, 35]. These methods require the manual annotation of many training images and retraining for each new task. Other methods use annotations to perform online learning, using user corrections as a source of supervision to update the model weights at test time [113, 5, 4, 121, 60, 140]. In contrast, MultiverSeg is trained only once on a large corpus of datasets once, and then can be used to segment new datasets at inference time without any retraining, using graphical interactions and previous segmentations as input.
@@ 3 Methods
Figure 2: MultiverSeg Architecture. The MultiverSeg network (left) takes as input a stack of target image inputs $q_{i}$ and a context set of image-segmentation pairs $\{(x_{l},y_{l})\}_{l=1}^{m}$. The target image inputs include a target image $x_{i}$, optional user interactions $u_{i,j}$, and a previous predicted segmentation $\hat{y}_{i,j-1}$ if available. The architecture is comprised of an encoder-decoder structure similar to a UNet [105]. We use a CrossBlock [16] (right) with additional normalization layers [6] to interact the features of the target image inputs $q_{i}$ with the features of the context set inputs $V=\{v_{l}\}_{l=1}^{m}$ throughout the network.
@@ 3.1 Problem Setup
We want to maximize the quality of predicted segmentations $\{\hat{y}_{i,k_{i}}^{t}\}_{i=1}^{N}$ for an entire dataset,
$\min\sum_{i=1}^{N}\mathcal{L}_{seg}(y_{i}^{t},\hat{y}_{i,k_{i}}^{t})$ (1)
while minimizing the total number of user interactions
$\sum_{i=1}^{N}\sum_{j=1}^{k_{i}}u_{i,j}^{t},$ (2)
where $\mathcal{L}_{seg}$ is a segmentation distance metric, $\{y_{i}^{t}\}_{i=1}^{N}$ are the ground truth segmentation maps, and $k_{i}$ is the number of steps of interaction for image $i$.
@@ 3.2 MultiverSeg
First image. When segmenting the first image $x_{0}^{t}$ of a new task $t$, the context set $S_{0}^{t}$ is empty. We build on learning-based interactive segmentation approaches [59, 126], to learn a function $g_{\phi}(x_{i}^{t};u_{i},\hat{y}^{t}_{i,j-1})$ that at step $j$ produces a segmentation $\hat{y}_{i,j}$ of image $x^{t}$, given a set of user interactions $u_{i,j}$ and a previously predicted segmentation $\hat{y}^{t}_{i,j-1}$. The interactions $u_{i}$, which may include positive or negative scribbles, positive or negative clicks, and bounding boxes, are provided by a user who has access to the image $x^{t}$ and previous prediction $\hat{y}_{i-1}^{t}$. For $g_{\phi}(\cdot)$, we use the pre-trained ScribblePrompt-UNet [126].
Subsequent images. For subsequent images $x_{i>0}^{t}$, the context set $S_{i}^{t}=\{(x_{l}^{t},\hat{y}_{l,k_{l}}^{t})\}_{l=0}^{i-1}$ encompass the previously segmented images and the resulting segmentation maps. We learn a function $f_{\theta}(x^{t};u_{i},\hat{y}^{t}_{i-1};S^{t})$ with parameters $\theta$ that leverages the set of user interactions $u_{i}$, previous prediction $\hat{y}^{t}_{i-1}$, and context set $S^{t}_{i}$ to produce a segmentation $\hat{y}_{i}$. As more images are segmented, the context set $S_{i}^{t}$ grows, leading to fewer interactions $u_{i}$ necessary for accurate segmentation of each subsequent image.
@@ 3.2.1 Architecture
Target image inputs. The target image inputs consist of the target image $x_{i}$ and graphical user interactions $u_{i,j}$, and a previous prediction $\hat{y}_{i,j-1}$ if available. The interactions may include bounding boxes, positive clicks and scribbles, and negative clicks and scribbles, represented as three intensity-based masks [126]. We stack the target inputs leading to five channels, where the first channel contains the input image, and the other channels contain the optional interactions and an optional previous prediction. When there are no interactions or there is no previous prediction, these channels are set to zero.
@@
We first use a cross-convolution layer to interact a target feature map $q$ with a set of context feature maps $V=\{v_{i}\}_{i=1}^{n}$:
@@
$\text{CrossBlock}(q,V;\ \theta_{z,q,v})=(q^{\prime},V^{\prime}),\text{where:}$: (4)
@@ 3.3 Training
We summarize the training process in Algorithm 1. During training, we first sample a random task $t$, and then sample a training example $(x_{i}^{t},y_{i}^{t})$ and context set $S_{i}^{t}=\{(x_{l}^{t},y_{l}^{t})\}_{l=0}^{n}$ of random size $n\in[0,N]$. We use ground truth segmentation labels in the context set.
We minimize the difference between the true segmentation $y^{t}$ and each of the $k$ iterative predictions $\hat{y}_{i,0},\dots,\hat{y}_{i,k}$ given a context set $S^{t}$,
@@
     $t\sim\mathcal{T}$triangleright Sample Task
     $(x_{i}^{t},y_{i}^{t})\sim t$triangleright Sample Target
     $n\sim U[0,N]$triangleright Sample Context Size
     $S^{t}\leftarrow\{(x_{l}^{t},y_{l}^{t})\}_{l\neq i}^{n}$triangleright Sample Context
     $x_{i}^{t},y_{i}^{t}\leftarrow\text{Aug}_{t}(x_{i}^{t},y_{i}^{t})$triangleright Augment Target
     $S^{t}\leftarrow\{\text{Aug}_{t}(x_{l}^{t},y_{l}^{t})\}_{l}^{n}$triangleright Augment Context
     $x_{i}^{t},y_{i}^{t},S^{t}\leftarrow\text{Aug}_{T}(x_{i}^{t},y_{i}^{t},S^{t})$triangleright Task Aug
@@
[…]
         $u_{i,j}^{t}\leftarrow h_{\psi}(y_{i}^{t},\hat{y}_{j-1})$triangleright Simulate Interactions
         $\hat{y}_{i,j}\leftarrow f_{\theta}(x_{i}^{t},u_{i,j}^{t},\hat{y}_{i,j-1};S^{t})$triangleright Predict Seg.
         $\ell_{j}\leftarrow\mathcal{L}_{\text{seg}}({y}_{i}^{t},\hat{y}_{i,j})$triangleright Compute Loss
@@
     $\theta\leftarrow\theta-\eta\nabla_{\theta}\sum_{j}\ell$triangleright Gradient Step
@@
Prompt Simulation. We simulate random combinations of scribbles, clicks and bounding boxes during training following prompt simulations procedures described in [126]. We simulate $k$ steps of interactive segmentation for each example during training. For the first step ($i=1$), we sample the combination of interactions (bounding box, clicks, scribbles) and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The initial interactions $u_{1}$ are simulated using the ground truth label $y^{t}$. In subsequent steps, we sample correction scribbles or clicks from the error region $\varepsilon_{i-1}^{t}$ between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth $y^{t}$. Since a user can make multiple corrections in each step, we sample $n_{cor}\sim U[n_{min},n_{max}]$ corrections (scribbles or clicks) per step.
@@ 4 Data
Task Diversity. We use a collection of 78 biomedical imaging datasets (Appendix B) and synthetically generated images and tasks [16, 101, 126]. The collection includes a diverse array of biomedical domains, such as eyes [44, 69, 86, 99, 116], thorax [108, 111, 112, 109, 100], spine [141, 80, 109], cells [142, 79, 17, 32, 18, 30, 38], skin [19], abdominal [12, 14, 41, 52, 55, 64, 66, 68, 74, 81, 100, 112, 87, 115, 104], neck [61, 98, 94, 97], brain [7, 33, 42, 62, 63, 88, 89, 91, 112, 2], bones [109, 40, 118], teeth [1, 50] and lesions [3, 137, 139, 112].
Task Definition. We define a 2D segmentation task as a combination of dataset, modality, axis (for 3D modalities), and binary label. For datasets with multiple segmentation labels, we consider each label separately as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice from each volume.
Data Augmentation. We perform both task augmentation and within-task data augmentation to increase the diversity of segmentation tasks [16]. For task augmentation, the same augmentation is applied to the target example and the entries of the context set to change the segmentation task. For within-task augmentation, we apply data augmentation where the parameters are randomly sampled for each target example and context set entry, to vary the examples within a task. Augmentations are applied prior to simulating the user interactions. We detail the augmentations in Sec. B.3.
Synthetic Data. Synthetic data can help improve generalization [16, 126, 34, 13]. We use some fully synthetic data (images and labels) similar to strategies used for in-context learning [16].
@@
To generate a set of $m+1$ images representing the same task, we duplicate $(x_{0},y_{synth})$, $m+1$ times and apply aggressive augmentations to vary the images and segmentation labels [138, 16]. We detail these augmentations and provide examples in Sec. B.2.

### v2

@@ Abstract
Medical researchers and clinicians often need to perform novel segmentation tasks on a set of related images. Existing methods for segmenting a new dataset are either interactive, requiring substantial human effort for each image, or require an existing set of previously labeled images.
@@
We demonstrate that MultiverSeg enables users to interactively segment new datasets efficiently, by amortizing the number of interactions per image to achieve an accurate segmentation. Compared to using a state-of-the-art interactive segmentation method, MultiverSeg reduced the total number of clicks by 36% and scribble steps by 25% to achieve 90% Dice on sets of images from unseen tasks. We release code and model weights at https://multiverseg.csail.mit.edu.
Figure 1: MultiverSeg enables users to rapidly segment new datasets. The MultiverSeg network takes as input an image to segment, user interactions, and a context set of previously segmented image-segmentation pairs (left). As the user completes more segmentations, those images and segmentations become additional inputs to the model, populating the context set. As the context set of labeled images grows, the number of interactions required to achieve an accurate segmentation decreases (right).
@@ 1 Introduction
Segmentation is an important step in biomedical image analysis pipelines. Biomedical and clinical researchers often acquire novel images types or identify new regions of interest, and need to perform new segmentation tasks. Typically, scientists want to segment the same region of interest in many similar images from a new dataset.
Manually segmenting images is labor-intensive and requires domain expertise. Interactive segmentation systems, in which a user provides a few clicks or scribbles on an image to produce a predicted segmentation, help to speed up the annotation of individual images. But with existing interactive segmentation systems, the user must independently repeat the same process for each image [61, 130, 88, 85, 19, 137]. Ideally, a system should be able to learn from experience, becoming more accurate as the user completes more segmentations from the same task.
We propose MultiverSeg, a new interactive system that, as more images are segmented, progressively reduces the number of user interactions needed to predict accurate segmentations (Fig. 1). MultiverSeg takes as input user interactions for a new image, along with a context set, of example (previously segmented) image-segmentation pairs. To segment a new dataset, the user begins by interactively segmenting the first image. Once completed, the example becomes an input to MultiverSeg, providing context for the segmentation of subsequent examples. The user interactively segments the next image with MultiverSeg using bounding boxes, clicks or scribbles. As the user labels more images, the context set grows and the number of interactions required to achieve the desired segmentation of subsequent images decreases, often to zero. Unlike existing interactive segmentations systems [130, 61, 85, 88, 140, 127, 78, 79, 135, 111, 126] where the work required to segment a dataset is linear in the number of images, MultiverSeg enables users to rapidly segment entire datasets.
@@ 2 Related Work
Interactive Segmentation. Recent interactive segmentation models can generalize to new segmentation tasks in medical [130, 88, 136] and natural images [61, 106]. While these methods are effective for segmenting single images, they require prohibitively extensive human interaction when segmenting large datasets. To incorporate new information, they must be retrained or fine-tuned. In contrast, at inference time MultiverSeg can be conditioned on example segmentations from the new task, dramatically reducing the human effort required for accurate segmentation.
@@
A few works have explored in-context segmentation using a context set of example images with user annotations on those example images. OnePrompt [132] segments a medical image given exactly one context example with click, scribble, bounding box or mask annotation. In contrast, MultiverSeg has mechanisms to facilitate a variable number of context set entries and enables users to incorporate interactions on the target image along with the context of previous segmentations, enabling substantially richer use cases. LabelAnything [25] is a few-shot framework that enables multi-label segmentation of a target natural image given a small context set of example images with click, box, or mask annotations on the examples. In contrast, MultiverSeg can leverage large context sets, and enables users to provide interactions on the target image and corrections to refine the prediction to get the desired result.
@@
Annotation-Efficient Learning. Another approach to segmenting a new dataset is to collect sparse annotations on many images and train an automatic segmentation model using these annotations for supervision. The annotations can be bounding boxes [67, 121], clicks [77, 81, 110] or scribbles [73, 72, 139, 86, 37]. These methods require the manual annotation of many training images and retraining for each new task. Other methods use annotations to perform online learning, using user corrections as a source of supervision to update the model weights at test time [117, 5, 4, 125, 62, 144]. In contrast, MultiverSeg is trained only once on a large corpus of datasets, and then can be used to segment new datasets at inference time without any retraining, using graphical interactions and previous segmentations as input.
@@ 3 Methods
Figure 2: MultiverSeg Architecture. The MultiverSeg network (left) takes as input a stack of target image inputs $q_{i}$ and a context set of image-segmentation pairs $\{(x_{l},y_{l})\}_{l=1}^{m}$. The target image inputs include a target image $x_{i}$, optional user interactions $u_{i,j}$, and a previous predicted segmentation $\hat{y}_{i,j-1}$, if available. The architecture is similar to a UNet [109]. However, we use a CrossBlock [16] (right) with additional normalization layers [6] to interact the features of the target image inputs $q_{i}$ with the features of the context set inputs $V=\{v_{l}\}_{l=1}^{m}$ throughout the network.
@@ 3.1 Problem Setup
We want to maximize the quality of predicted segmentations $\{\hat{y}_{i,k_{i}}^{t}\}_{i=1}^{N}$ for an entire dataset, $\min\sum_{i=1}^{N}\mathcal{L}_{seg}(y_{i}^{t},\hat{y}_{i,k_{i}}^{t})$ while minimizing the total number of user interactions $\sum_{i=1}^{N}\sum_{j=1}^{k_{i}}u_{i,j}^{t}$, where $\mathcal{L}_{seg}$ is a segmentation distance metric, $\{y_{i}^{t}\}_{i=1}^{N}$ are the ground truth segmentation maps, and $k_{i}$ is the number of steps of interaction for image $i$.
@@ 3.2 MultiverSeg
First image. When segmenting the first image $x_{0}^{t}$ of a new task $t$, the context set $S_{0}^{t}$ is empty. We build on learning-based interactive segmentation approaches [61, 130], to learn a function $g_{\phi}(x_{i}^{t};u_{i},\hat{y}^{t}_{i,j-1})$ that at step $j$ produces a segmentation $\hat{y}_{i,j}$ of image $x^{t}$, given a set of user interactions $u_{i,j}$ and a previously predicted segmentation $\hat{y}^{t}_{i,j-1}$. The interactions $u_{i}$, which may include positive or negative scribbles, positive or negative clicks, and bounding boxes, are provided by a user who has access to the image $x^{t}$ and previous prediction $\hat{y}_{i-1}^{t}$. For $g_{\phi}(\cdot)$, we use the pre-trained ScribblePrompt-UNet model [130].
Subsequent images. For subsequent images $x_{i>0}^{t}$, the context set $S_{i}^{t}=\{(x_{l}^{t},\hat{y}_{l,k_{l}}^{t})\}_{l=0}^{i-1}$ encompass the previously segmented images and the resulting segmentation maps. We learn a function $f_{\theta}(x^{t};u_{i},\hat{y}^{t}_{i-1};S^{t})$ with parameters $\theta$ that leverages the set of user interactions $u_{i}$, previous prediction $\hat{y}^{t}_{i-1}$, and context set $S^{t}_{i}$ to produce a segmentation $\hat{y}_{i}$. As more images are segmented, the context set $S_{i}^{t}$ grows, leading to fewer interactions $u_{i}$ needed for accurate segmentation of each subsequent image.
@@ 3.2.1 Architecture
Target image inputs. The target image inputs consist of the target image $x_{i}$ and graphical user interactions $u_{i,j}$, and a previous prediction $\hat{y}_{i,j-1}$, if available. The interactions may include bounding boxes, positive clicks and scribbles, and negative clicks and scribbles, represented as three intensity-based masks [130]. We stack the target inputs leading to five channels, where the first channel contains the input image, and the other channels contain the optional interactions and an optional previous prediction. When there are no interactions or there is no previous prediction, these channels are set to zero.
@@
We use a cross-convolution layer to interact a target feature map $q$ with a set of context feature maps $V=\{v_{i}\}_{i=1}^{n}$:
@@
$\text{CrossBlock}(q,V;\ \theta_{z,q,v})=(q^{\prime},V^{\prime}),\text{where:}$ (2)
@@ 3.3 Training
We summarize the training process in Algorithm 1. During training, we first sample a random task $t$, and then sample a training example $(x_{i}^{t},y_{i}^{t})$ and context set $S_{i}^{t}=\{(x_{l}^{t},y_{l}^{t})\}_{l=0}^{n}$ of random size $n\in[0,N]$. We use ground truth segmentation labels in the context set during training.
We minimize the difference between the true segmentation $y^{t}$ and each of the $k$ iterative predictions $\hat{y}_{i,0},\dots,\hat{y}_{i,k}$, given a context set $S^{t}$,
@@
  $t\sim\mathcal{T}$ $\triangleright$ Sample Task
  $(x_{i}^{t},y_{i}^{t})\sim t$ $\triangleright$ Sample Target
  $n\sim U[0,N]$ $\triangleright$ Sample Context Size
  $S^{t}\leftarrow\{(x_{l}^{t},y_{l}^{t})\}_{l\neq i}^{n}$ $\triangleright$ Sample Context
  $x_{i}^{t},y_{i}^{t}\leftarrow\text{Aug}_{t}(x_{i}^{t},y_{i}^{t})$ $\triangleright$ Augment Target
  $S^{t}\leftarrow\{\text{Aug}_{t}(x_{l}^{t},y_{l}^{t})\}_{l}^{n}$ $\triangleright$ Augment Context
  $x_{i}^{t},y_{i}^{t},S^{t}\leftarrow\text{Aug}_{T}(x_{i}^{t},y_{i}^{t},S^{t})$ $\triangleright$ Task Aug
@@
  for $j=1,\ldots,\text{NumInteractionSteps}$ do
   $u_{i,j}^{t}\leftarrow h_{\psi}(y_{i}^{t},\hat{y}_{j-1})$ $\triangleright$ Simulate Interactions
   $\hat{y}_{i,j}\leftarrow f_{\theta}(x_{i}^{t},u_{i,j}^{t},\hat{y}_{i,j-1};S^{t})$ $\triangleright$ Predict Seg.
   $\ell_{j}\leftarrow\mathcal{L}_{\text{seg}}({y}_{i}^{t},\hat{y}_{i,j})$ $\triangleright$ Compute Loss
@@
  $\theta\leftarrow\theta-\eta\nabla_{\theta}\sum_{j}\ell$ $\triangleright$ Gradient Step
@@
Prompt Simulation. We simulate random combinations of scribbles, clicks and bounding boxes during training following the prompt simulation procedures described in [130]. We simulate $k$ steps of interactive segmentation for each example during training. For the first step ($i=1$), we sample the combination of interactions (bounding box, clicks, scribbles) and the number of initial positive and negative interactions $n_{pos},n_{neg}\sim U[n_{min},n_{max}]$. The initial interactions $u_{1}$ are simulated using the ground truth label $y^{t}$. In subsequent steps, we sample correction scribbles or clicks from the error region $\varepsilon_{i-1}^{t}$ between the last prediction $\hat{y}^{t}_{i-1}$ and the ground truth $y^{t}$. Since a user can make multiple corrections in each step, we sample $n_{cor}\sim U[n_{min},n_{max}]$ corrections (scribbles or clicks) per step.
@@ 4 Data
Task Diversity. We use a collection of 79 biomedical imaging datasets (Appendix C) and synthetically generated images and tasks [16, 104, 130]. The collection includes a diverse array of biomedical domains, such as eyes [46, 71, 89, 102, 120], thorax [112, 115, 116, 113, 103], spine [145, 83, 113], cells [146, 82, 17, 34, 18, 32, 40], skin [21], abdomen [12, 14, 43, 54, 57, 66, 68, 70, 76, 84, 103, 116, 90, 119, 107], neck [63, 101, 97, 100], brain [7, 35, 44, 64, 65, 91, 92, 94, 116, 2], bones [113, 42, 122, 129], teeth [1, 52] and lesions [3, 141, 143, 116].
Task Definition. We define a 2D segmentation task as a combination of dataset, modality, axis (for 3D modalities), and binary label. For datasets with multiple segmentation labels, we consider each label as a binary segmentation task and for 3D modalities we use the slice with maximum label area and the middle slice from each volume.
Data Augmentation. We perform both task augmentation and within-task data augmentation to increase the diversity of segmentation tasks [16]. For task augmentation, the same augmentation is applied to the target example and the entries of the context set to change the segmentation task. For within-task augmentation, we apply data augmentation where the parameters are randomly sampled for each target example and context set entry, to vary the examples within a task. Augmentations are applied prior to simulating the user interactions. We detail the augmentations in Sec. C.3.
Synthetic Data. Synthetic data can help improve generalization [16, 130, 36, 13]. We use fully synthetic data (images and labels) similar to strategies used for in-context learning [16].
@@
To generate a set of $m+1$ images representing the same task, we duplicate $(x_{0},y_{synth})$, $m+1$ times and apply aggressive augmentations to vary the images and segmentation labels [142, 16]. We detail these augmentations and provide examples in Sec. C.2.